What is the answer to (5c-1)(5c+1)
Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

(a-b) * (a+b) = a^2 - b^2

by that rule, (5c-1)*(5x+1) should be equal to (5c)^2 - (1)^2 which is equal to 5^2c^2 - 1^2 which is equal to 25c^2 - 1.

let's see if that works.

using the distributive law of multiplication:

(5c-1)*(5c+1) = 5c*5c + 1*5c - 1*5c - 1*1

the + 1*5x and the - 1*5x cancel out and you are left with:

5c*5c - 1*1 which is equal to (5c)^2 - 1^2 which is equal to 5^2*c^2 - 1 which is equal to 25c^2 - 1.

the rule works.

if you recognize the pattern, you can apply the rule.

if you don't, then you just multiply the factors as you would normally and you will get the same result.




Answer by Alan3354(69443)

Depends on the question.